# Synthetic Augmentation in Imbalanced Learning: When It Helps, When It Hurts, and How Much to Add

### v1

@@ Abstract
Imbalanced classification, where one class is observed far less frequently than the other, often causes standard training procedures to prioritize the majority class and perform poorly on rare but important cases. A classic and widely used remedy is to augment the minority class with synthetic examples, but two basic questions remain under-resolved: when does synthetic augmentation actually help, and how many synthetic samples should be generated?
We develop a unified statistical framework for synthetic augmentation in imbalanced learning, studying models trained on imbalanced data augmented with synthetic minority samples and evaluated under the balanced population risk. Our theory shows that synthetic data is not always beneficial. In a “local symmetry” regime, imbalance is not the dominant source of error near the balanced optimum, so adding synthetic samples cannot improve learning rates and can even degrade performance by amplifying generator mismatch. When augmentation can help (a “local asymmetry” regime), the optimal synthetic size depends on generator accuracy and on whether the generator’s residual mismatch is directionally aligned with the intrinsic majority–minority shift. This structure can make the best synthetic size deviate from naive full balancing, sometimes by a small refinement and sometimes substantially when generator bias is systematic. Practically, we recommend Validation-Tuned Synthetic Size (VTSS): select the synthetic size by minimizing balanced validation loss over a range centered near the fully balanced baseline, while allowing meaningful departures when the data indicate them. Simulations and a real sepsis prediction study support the theory and illustrate when synthetic augmentation helps, when it cannot, and how to tune its quantity effectively.
@@ 1 Introduction
Imbalanced classification, where one class is observed far less frequently than the other, is a pervasive obstacle in modern statistical learning. Standard empirical risk minimization tends to favor overall accuracy, which often leads to systematically degraded detection of the minority class. This failure mode is especially consequential when the rare events are clinically, economically, or operationally important, as in medical data analysis (Salmi et al., 2024; Hung et al., 2022; Faviez et al., 2020), finance (Chen et al., 2024b; Tong and Shen, 2023; Subudhi and Panigrahi, 2018), and industrial anomaly detection (Bougaham et al., 2024; Piciarelli et al., 2021). These challenges motivate a renewed statistical focus on mitigating imbalance in a principled way, without turning augmentation into a purely heuristic step (Chen et al., 2024a).
A classic response to imbalance is to augment the minority class with additional (real or synthetic) samples, so that model training is less dominated by the majority class. The simplest mechanism is bootstrap-style oversampling, i.e., resampling minority observations with replacement (Efron and Tibshirani, 1994). A widely used alternative is SMOTE, which creates new minority examples by interpolating between a minority observation and its neighbors (Chawla et al., 2002). SMOTE has inspired many extensions, including ADASYN (He et al., 2008), Borderline-SMOTE (Han et al., 2005), safe-level-SMOTE (Bunkhumpornpat et al., 2009), and DBSMOTE (Bunkhumpornpat et al., 2012). Related augmentation strategies include Mixup (Zhang et al., 2017) and conditional data synthesis approaches (Tian and Shen, 2025). Beyond classical generators, modern deep generative models can also serve as synthetic sample generators, including VAEs (Kingma and Welling, 2013), normalizing flows (Papamakarios et al., 2021), GANs (Goodfellow et al., 2014), and diffusion/score-based models (Ho et al., 2020; Song et al., 2020). More recently, attention-based large language models have also been explored for synthetic data generation (Nakada et al., 2024). For a broad view of deep generative synthetic data across modalities and domains, see Eigenschink et al. (2023).
@@
First, synthetic augmentation is not guaranteed to help. Synthetic samples are generated from an estimated minority distribution. Augmentation can reduce the impact of imbalance, but it can simultaneously introduce distributional mismatch (because the generator is imperfect) and alter the effective variance of learning. As a result, the net effect on the target performance criterion can be positive or negative depending on data geometry and generator quality.
Second, even when augmentation is beneficial, it is unclear how many synthetic samples should be added. A common practice is naive balancing: generate enough synthetic minority samples so that the minority-plus-synthetic count matches the majority count. This rule is simple and often reasonable, but it is largely heuristic. Crucially, synthetic samples generally do not carry the same information content as independent real samples, and their mismatch can interact with the sample size in nontrivial ways.
@@
This paper develops a unified framework that treats synthetic augmentation as a controllable statistical operation and studies the risk trade-offs it induces. Our focus is the performance under a balanced evaluation criterion (i.e., treating the two classes symmetrically at the population level), which is common in imbalanced learning when minority detection is a primary goal. Within this framework, augmentation is characterized by two ingredients: (1) the synthetic size (how many synthetic minority samples are added), and (2) the generator mismatch (how the synthetic minority distribution differs from the true minority distribution). Our analysis shows that these two factors jointly determine whether augmentation helps, how it helps, and when it can be harmful.
@@
Augmenting the minority class is often beneficial, but not universal. The key distinction is whether imbalance creates a first-order distortion of the learning objective near the population optimum under the balanced criterion. In many problems, the majority and minority contribute differently in the optimization-relevant directions; in this case, changing the effective class proportions can meaningfully shift the learned solution toward the balanced target. We refer to this as a local asymmetry regime, and it is precisely the setting in which synthetic augmentation can help.
However, our analysis also identifies an important regime where synthetic augmentation cannot help, even under severe imbalance. In this regime, the learning problem exhibits a form of “local symmetry”: near the balanced optimum, the two classes already exert equal first-order influence in the directions that matter for optimization. Then imbalance is not the bottleneck, and adding synthetic minority samples provides little benefit. Moreover, if the generator is imperfect, adding more synthetic samples can amplify mismatch effects and degrade performance. We develop this phenomenon in detail in Section 5.
@@
Naive balancing is not universally optimal, although it is often a sensible default. Whether it is optimal depends on generator quality and, more subtly, on whether the residual synthetic mismatch aligns with the intrinsic majority–minority difference in the learning problem. When the generator is consistent and there is no usable directional structure in its remaining mismatch, naive balancing is essentially rate-optimal: it corrects the leading imbalance effect, and finer tuning yields at most constant-factor improvements. In contrast, when the mismatch is directionally aligned with the majority–minority discrepancy, then a small adjustment around naive balancing can reduce the leading bias and yield a measurable improvement. Finally, for systematically biased (inconsistent) generators, tuning the synthetic size may be necessary for consistency, and the best choice can be far from naive balancing when bias cancellation is possible; when the mismatch direction is not aligned, adjusting the quantity alone cannot remove the bias.
@@
In real applications, the relevant regime and mismatch direction are rarely known a priori. Motivated by the theory, we propose a simple and robust procedure, Validation-Tuned Synthetic Size (VTSS): sweep a range of synthetic sizes (typically near naive balancing) and choose the value that minimizes balanced validation loss. VTSS is implementation-friendly, captures potential gains from nontrivial sizing, and automatically avoids harmful over-synthesis in regimes where synthetic data does not help.
The main messages of our theory are summarized in Figure 1, which maps the local asymmetry/symmetry regimes and generator quality to when synthetic augmentation helps and how the synthetic size should be chosen.
Figure 1: Flowchart of main results.
@@
Synthetic minority augmentation has a long history in statistics and machine learning, including bootstrap-style oversampling (Efron and Tibshirani, 1994), SMOTE (Chawla et al., 2002) and its variants such as Borderline-SMOTE (Han et al., 2005), ADASYN (He et al., 2008), safe-level-SMOTE (Bunkhumpornpat et al., 2009), and DBSMOTE (Bunkhumpornpat et al., 2012), as well as related augmentation strategies such as Mixup (Zhang et al., 2017) and conditional synthesis (Tian and Shen, 2025). Deep generative approaches include VAEs (Kingma and Welling, 2013), normalizing flows (Papamakarios et al., 2021), GANs (Goodfellow et al., 2014), and diffusion/score-based models (Ho et al., 2020; Song et al., 2020); LLM-based oversampling has also been studied (Nakada et al., 2024). A broader survey of deep generative approaches to synthetic data can be found in Eigenschink et al. (2023).
Recent theoretical work has begun to clarify statistical properties of learning with synthetic data. Nakada et al. (2024) provides theoretical guarantees for when LLM-generated minority samples can improve learning under imbalance and discusses interactions between generator quality and augmentation size. Lyu et al. (2025) emphasizes that naively treating synthetic samples as real can introduce systematic bias and proposes bias-correction strategies with extensions beyond standard imbalanced classification. For uncertainty quantification and inference, Keret and Shojaie (2025) studies inference with AI-generated synthetic data for generalized linear models, while Räisä et al. (2025) develops a Bayesian framework that accounts for the synthetic generation mechanism and establishes conditions for consistent inference. Xia and Klusowski (2026) frames imbalanced classification as transfer learning under label shift and highlights settings in which oversampling incurs a cost, including curse-of-dimensionality effects for SMOTE. Ahmad et al. (2025) develops uniform concentration bounds comparing empirical risk on synthetic samples to the true minority risk and derives excess-risk guarantees for kernel-based classifiers.
@@
The rest of this paper is organized as follows. Section 2 introduces the notation and formal setup for imbalanced classification with synthetic minority augmentation, and defines the balanced evaluation objective used throughout. Section 3 develops our main excess-risk characterization, which makes explicit how performance depends on both class-proportion imbalance and generator-induced distributional mismatch. Section 4 studies the local asymmetry regime in which synthetic augmentation can improve learning, and shows how the optimal synthetic size depends on generator quality and the directional structure of its bias. Section 5 identifies a local symmetry regime in which imbalance is not the dominant bottleneck, so synthetic augmentation cannot improve rates and may degrade performance when the generator is imperfect. Section 6 then presents our practical procedure, Validation-Tuned Synthetic Size (VTSS), which selects the synthetic size by minimizing balanced validation loss. Section 7 provides simulation studies that verify the qualitative predictions of the theory and evaluate VTSS. Section 8 concludes with discussion and future directions. Proofs and technical details are deferred to the supplementary materials.
@@ 2 Notation and Problem Formulation
For nonnegative sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}\lesssim b_{n}$ (resp. $a_{n}\gtrsim b_{n}$) if there exists a universal constant $C>0$ such that $a_{n}\leq Cb_{n}$ (resp. $a_{n}\geq Cb_{n}$). We write $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold, and write $a_{n}\ll b_{n}$ (resp. $a_{n}\gg b_{n}$) if $a_{n}/b_{n}\to 0$ (resp. $a_{n}/b_{n}\to\infty$).
Vectors are denoted by boldface letters (e.g., $\bm{x},\bm{y},\bm{u},\bm{v}$), and $\|\cdot\|$ denotes the Euclidean norm. Gradients and Hessians with respect to $\bm{\theta}$ are denoted by $\nabla$ and $\nabla^{2}$. For two nonzero vectors $\bm{u},\bm{v}$ of the same dimension, we denote their angle and the sine value by
$\angle(\bm{u},\bm{v}):=\arccos\left(\frac{\bm{u}^{\top}\bm{v}}{\|\bm{u}\|\|\bm{v}\|}\right),~~\sin\angle(\bm{u},\bm{v})=\sqrt{1-\left(\frac{\bm{u}^{\top}\bm{v}}{\|\bm{u}\|\|\bm{v}\|}\right)^{2}}.$
For a symmetric matrix $A$, let $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denote its largest and smallest eigenvalues. For symmetric matrices $A$ and $B$, we write $A\succeq B$ (resp. $A\succ B$) if ABA-B is positive semidefinite (resp. positive definite).
Throughout, we work in an asymptotic regime where the data-generating distributions are fixed while the sample size grows. We use standard stochastic order notation $O_{P}(\cdot)$, $o_{P}(\cdot)$, $\Omega_{P}(\cdot)$, and $\Theta_{P}(\cdot)$. In particular, for sequences indexed by $(n_{0},n_{1})$, we write $x_{(n_{0},n_{1})}=O_{P}(a_{(n_{0},n_{1})})$ as $n_{0},n_{1}\to\infty$ if $x_{(n_{0},n_{1})}/a_{(n_{0},n_{1})}$ is stochastically bounded, and write $x_{(n_{0},n_{1})}=o_{P}(1)$ if $x_{(n_{0},n_{1})}\to 0$ in probability. We write $x_{(n_{0},n_{1})}=\Omega_{P}(a_{(n_{0},n_{1})})$ if $x_{(n_{0},n_{1})}/a_{(n_{0},n_{1})}$ is stochastically bounded away from zero, and $x_{(n_{0},n_{1})}=\Theta_{P}(a_{(n_{0},n_{1})})$ if both $O_{P}(\cdot)$ and $\Omega_{P}(\cdot)$ hold.
@@
We consider binary classification under class imbalance, where $y=0$ denotes the majority class and $y=1$ denotes the minority class. Let the class-conditional distributions be
$\bm{x}\mid(y=0)\sim\mathcal{P}_{0},\qquad\bm{x}\mid(y=1)\sim\mathcal{P}_{1}.$
We observe $n_{0}$ majority samples and $n_{1}$ minority samples, and we may augment the data with $\tilde{n}$ synthetic minority samples. Synthetic covariates are drawn from a generator-induced distribution $\mathcal{P}_{\text{syn}}$; since the generator is typically trained on the observed data, $\mathcal{P}_{\text{syn}}$ may depend on $(n_{0},n_{1})$ (more precisely, $\mathcal{P}_{\text{syn},n_{0},n_{1}}$), though we suppress this dependence for notational simplicity.
Let $\ell(\bm{\theta};\bm{x},y)$ be a loss function with parameter $\bm{\theta}\in\Theta$. Common choices include cross-entropy/logistic loss, squared loss, hinge loss, and exponential loss. To treat the two classes symmetrically at the population level, we define the balanced population risk
@@
and denote its minimizer by
$\bm{\theta}^{*}=\mathop{\rm arg\min}_{\bm{\theta}\in\Theta}\mathcal{R}(\bm{\theta}).$
Augmenting the training data with $\tilde{n}$ synthetic minority samples changes the effective class proportions. Let
$\pi_{0}=\frac{n_{0}}{n_{0}+n_{1}+\tilde{n}},\qquad\pi_{1}=\frac{n_{1}}{n_{0}+n_{1}+\tilde{n}},\qquad\tilde{\pi}=\frac{\tilde{n}}{n_{0}+n_{1}+\tilde{n}}$,
so that $\pi_{0}+\pi_{1}+\tilde{\pi}=1$. The resulting synthetic population risk is
@@
with minimizer $\tilde{\bm{\theta}}=\mathop{\rm arg\min}_{\bm{\theta}\in\Theta}\tilde{\mathcal{R}}(\bm{\theta})$.
In practice, we observe $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{0}+n_{1}}$ and generate synthetic samples
$\{\tilde{\bm{x}}_{i}\}_{i=1}^{\tilde{n}}\sim\mathcal{P}_{\text{syn}},\qquad\tilde{y}_{i}\equiv 1.$
We then minimize the empirical augmented risk
@@
where $\{\bm{x}_{i}^{(0)}\}_{i=1}^{n_{0}}$ and $\{\bm{x}_{i}^{(1)}\}_{i=1}^{n_{1}}$ are the observed majority and minority samples. Let
$\hat{\bm{\theta}}=\mathop{\rm arg\min}_{\bm{\theta}\in\Theta}\hat{\mathcal{R}}(\bm{\theta}).$
Our goal is to estimate the balanced-risk optimizer $\bm{\theta}^{*}$ from imbalanced data, and to understand how synthetic augmentation, through $\tilde{n}$ and $\mathcal{P}_{\text{syn}}$, affects the balanced excess risk $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$.

### v2

@@ Abstract
Imbalanced classification often causes standard training procedures to prioritize the majority class and perform poorly on rare but important cases. A classic and widely used remedy is to augment the minority class with synthetic samples, but two basic questions remain under-resolved: when does synthetic augmentation actually help, and how many synthetic samples should be generated?
We develop a unified statistical framework for synthetic augmentation in imbalanced learning, studying models trained on imbalanced data augmented with synthetic minority samples. Our theory shows that synthetic data is not always beneficial. In a “local symmetry” regime, imbalance is not the dominant source of error, so adding synthetic samples cannot improve learning rates and can even degrade performance by amplifying generator mismatch. When augmentation can help (“local asymmetry”), the optimal synthetic size depends on generator accuracy and on whether the generator’s residual mismatch is directionally aligned with the intrinsic majority–minority shift. This structure can make the best synthetic size deviate from naive full balancing. Practically, we recommend Validation-Tuned Synthetic Size (VTSS): select the synthetic size by minimizing balanced validation loss over a range centered near the fully balanced baseline, while allowing meaningful departures. Extensive simulations and real data analysis further support our findings.
Abstract
In the supplementary materials, we discuss related classifiers and synthetic data generators and provide proofs of all technical results.
Keywords: Data Imbalance, Synthetic Oversampling, Synthetic Data Size.
@@ 1 Introduction
Imbalanced classification, where one class is observed far less frequently than the other, is a pervasive obstacle in modern statistical learning. Standard empirical risk minimization tends to favor overall accuracy, which often leads to systematically degraded detection of the minority class. This failure mode is especially consequential when the rare events are clinically, economically, or operationally important, as in medical data analysis (Salmi et al. 2024), finance (Chen, Calabrese & Martin-Barragan 2024), and industrial anomaly detection (Bougaham et al. 2024). These challenges motivate a renewed statistical focus on mitigating imbalance in a principled way, without turning augmentation into a purely heuristic step (Chen, Yang, Yu, Shi & Chen 2024).
A classic response to imbalance is to augment the minority class with additional (real or synthetic) samples, so that model training is less dominated by the majority class. The simplest mechanism is bootstrap-style oversampling, i.e., resampling minority observations with replacement (Efron & Tibshirani 1994). A widely used alternative is SMOTE, which creates new minority examples by interpolating between a minority observation and its neighbors (Chawla et al. 2002). SMOTE has inspired many extensions, including ADASYN (He et al. 2008), Borderline-SMOTE (Han et al. 2005), safe-level-SMOTE (Bunkhumpornpat et al. 2009), and DBSMOTE (Bunkhumpornpat et al. 2012). Related augmentation strategies include Mixup (Zhang et al. 2017) and conditional data synthesis approaches (Tian & Shen 2025). Beyond classical generators, modern deep generative models can also serve as synthetic sample generators, including VAEs (Kingma & Welling 2013), normalizing flows (Papamakarios et al. 2021), GANs (Goodfellow et al. 2014), and diffusion/score-based models (Ho et al. 2020, Song, Sohl-Dickstein, Kingma, Kumar, Ermon & Poole 2020). More recently, attention-based large language models have also been explored for synthetic data generation (Nakada et al. 2024). For a broad view of deep generative synthetic data across modalities and domains, see Eigenschink et al. (2023).
@@
First, synthetic augmentation is not guaranteed to help. Because synthetic samples are generated from an estimated minority distribution, augmentation can reduce the impact of imbalance while also introducing generator mismatch and changing the effective variance. Its overall effect on the target performance criterion therefore depends on both the data geometry and the quality of the generator.
Second, even when augmentation is beneficial, it is unclear how many synthetic samples should be added. A common practice is naive balancing: generate enough synthetic minority samples so that the minority-plus-synthetic count matches the majority count. This rule is simple and often reasonable, but it is largely heuristic. Synthetic samples generally do not carry the same information as independent real samples, and their mismatch can interact with sample size in nontrivial ways.
@@
This paper develops a unified framework that treats synthetic augmentation as a controllable statistical operation and analyzes the risk trade-offs it induces under a balanced evaluation criterion, i.e., one that treats the two classes symmetrically at the population level. This criterion is common in imbalanced learning when minority detection is a primary goal. Our framework applies to a broad class of training losses, classification rules, and synthetic-sample generators, and it is parameterized by two quantities: the synthetic size and the generator mismatch. Our results show how these two factors jointly determine the statistical value of augmentation.
@@
Not always. The key distinction is whether imbalance creates a first-order distortion of the learning objective near the population optimum under the balanced criterion. When the two classes exert different first-order influence in optimization-relevant directions, changing the effective class proportions can shift the learned solution toward the balanced target. We call this the local asymmetry regime, and it is precisely the setting in which synthetic augmentation can improve performance.
Our analysis also identifies an important complementary regime in which synthetic augmentation cannot help, even under severe imbalance. In this local symmetry regime, the two classes already exert equal first-order influence near the balanced optimum in the directions that matter for optimization. Then imbalance is not the bottleneck, so adding synthetic minority samples yields little benefit; if the generator is imperfect, it can instead amplify mismatch effects and degrade performance. We develop this phenomenon in detail in Section 5.
@@
Naive balancing is often a sensible default, but it is not universally optimal. Its optimality depends on generator quality and on whether the residual synthetic mismatch aligns with the intrinsic majority–minority difference in the learning problem. When the generator is consistent and its remaining mismatch has no exploitable directional structure, naive balancing is essentially rate-optimal: it removes the leading imbalance effect, and finer tuning can improve performance only by a constant factor. By contrast, when the mismatch is directionally aligned with the majority–minority discrepancy, a small adjustment around naive balancing can further reduce the leading bias and produce a measurable gain. For systematically biased (inconsistent) generators, tuning the synthetic size may even be necessary for consistency, and the optimal choice can be far from naive balancing when bias cancelation is possible; when the mismatch direction is not aligned, changing the quantity alone cannot remove the bias.
@@
In real applications, the relevant regime and mismatch direction are rarely known a priori. Motivated by the theory, we propose a simple and robust procedure, Validation-Tuned Synthetic Size (VTSS): sweep a range of synthetic sizes and choose the value that minimizes balanced validation loss. VTSS is implementation-friendly, captures potential gains from nontrivial sizing, and automatically avoids harmful over-synthesis in regimes where synthetic data does not help.
The main messages of our theory are summarized in Figure 1, which maps the local asymmetry/symmetry regimes and generator quality to when synthetic augmentation is helpful and how the synthetic size should be chosen.
Figure 1: Flowchart of main results
Our framework accommodates a broad class of training losses, classification rules, and synthetic-sample generators. For completeness, the Supplementary Materials provide additional details and examples for the losses (Section 10), classifiers (Section 11), and generators (Section 12) considered in this work.
@@
Synthetic minority augmentation has a long history in statistics and machine learning, including bootstrap-style oversampling (Efron & Tibshirani 1994), SMOTE (Chawla et al. 2002) and its variants such as Borderline-SMOTE (Han et al. 2005), ADASYN (He et al. 2008), safe-level-SMOTE (Bunkhumpornpat et al. 2009), and DBSMOTE (Bunkhumpornpat et al. 2012), as well as related augmentation strategies such as Mixup (Zhang et al. 2017) and conditional synthesis (Tian & Shen 2025). Deep generative approaches include VAEs (Kingma & Welling 2013), normalizing flows (Papamakarios et al. 2021), GANs (Goodfellow et al. 2014), and diffusion/score-based models (Ho et al. 2020, Song, Sohl-Dickstein, Kingma, Kumar, Ermon & Poole 2020); LLM-based oversampling has also been studied (Nakada et al. 2024). A broader survey of deep generative approaches to synthetic data can be found in Eigenschink et al. (2023).
Relatedly, class imbalance can also be addressed without generating data, through the choice of objective, decision rule, and evaluation metric. Existing work considered three common paradigms: (i) classical risk minimization (often paired with reweighting or resampling), (ii) cost-sensitive learning that minimizes a weighted combination of type I and type II errors, and (iii) the Neyman–Pearson (NP) paradigm, which controls type I error at a prespecified level while minimizing type II error (Rigollet & Tong 2011, Tong et al. 2020, Feng et al. 2021). These perspectives motivate alternatives to oversampling such as loss reweighting, threshold adjustment, and constraint-based formulations.
Recent theoretical work has begun to clarify statistical properties of learning with synthetic data. Nakada et al. (2024) provides theoretical guarantees for when LLM-generated minority samples can improve learning under imbalance and discusses interactions between generator quality and augmentation size. Lyu et al. (2025) emphasizes that naively treating synthetic samples as real can introduce systematic bias and proposes bias-correction strategies with extensions beyond standard imbalanced classification. For uncertainty quantification and inference, Keret & Shojaie (2025) studies inference with AI-generated synthetic data for generalized linear models, while Räisä et al. (2025) develops a Bayesian framework that accounts for the synthetic generation mechanism and establishes conditions for consistent inference. Xia & Klusowski (2026) frames imbalanced classification as transfer learning under label shift and highlights settings in which oversampling incurs a cost, including curse-of-dimensionality effects for SMOTE. Ahmad et al. (2025) develops uniform concentration bounds comparing empirical risk on synthetic samples to the true minority risk and derives excess-risk guarantees for kernel-based classifiers.
@@
The rest of this paper is organized as follows. Section 2 introduces the notation and formal setup for imbalanced classification with synthetic minority augmentation, and defines the balanced evaluation objective used throughout. Section 3 develops our main excess-risk characterization, which makes explicit how performance depends on both class-proportion imbalance and generator-induced distributional mismatch. Section 4 studies the local asymmetry regime in which synthetic augmentation can improve learning, and shows how the optimal synthetic size depends on generator quality and the directional structure of its bias. Section 5 identifies a local symmetry regime in which imbalance is not the dominant bottleneck, so synthetic augmentation cannot improve rates and may degrade performance when the generator is imperfect. Section 6 then presents our practical procedure, Validation-Tuned Synthetic Size (VTSS), which selects the synthetic size by minimizing balanced validation loss. Section 7 provides simulation studies that verify the qualitative predictions of the theory and evaluate VTSS, and Section 8 reports a real-data application using electronic health records (EHR) data. Section 9 concludes with discussion and future directions. Proofs and additional technical details are deferred to the Supplementary Materials, which also include further examples for the losses (Section 10), classifiers (Section 11), and synthetic-sample generators (Section 12) covered by our framework.
@@ 2 Notation and Problem Formulation
For nonnegative sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}\lesssim b_{n}$ (resp. $a_{n}\gtrsim b_{n}$) if there exists a universal constant $C>0$ such that $a_{n}\leq Cb_{n}$ (resp. $a_{n}\geq Cb_{n}$). We write $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold, and write $a_{n}\ll b_{n}$ (resp. $a_{n}\gg b_{n}$) if $a_{n}/b_{n}\to 0$ (resp. $a_{n}/b_{n}\to\infty$). Vectors are denoted by boldface letters (e.g., $\bm{x},\bm{y},\bm{u},\bm{v}$), and $\|\cdot\|$ denotes the Euclidean norm. Gradients and Hessians with respect to $\bm{\theta}$ are denoted by $\nabla$ and $\nabla^{2}$. For two nonzero vectors $\bm{u},\bm{v}$ of the same dimension, we denote their angle and the sine value by $\angle(\bm{u},\bm{v}):=\arccos\left(\frac{\bm{u}^{\top}\bm{v}}{\|\bm{u}\|\|\bm{v}\|}\right),~\sin\angle(\bm{u},\bm{v})=\sqrt{1-\left(\frac{\bm{u}^{\top}\bm{v}}{\|\bm{u}\|\|\bm{v}\|}\right)^{2}}.$ For a symmetric matrix $A$, let $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denote its largest and smallest eigenvalues. For symmetric matrices $A$ and $B$, we write $A\succeq B$ (resp. $A\succ B$) if ABA-B is positive semidefinite (resp. positive definite). Throughout, we work in an asymptotic regime where the data-generating distributions are fixed while the sample size grows. We use standard stochastic order notation $O_{P}(\cdot)$, $o_{P}(\cdot)$, $\Omega_{P}(\cdot)$, and $\Theta_{P}(\cdot)$. In particular, for sequences indexed by $(n_{0},n_{1})$, we write $x_{(n_{0},n_{1})}=O_{P}(a_{(n_{0},n_{1})})$ as $n_{0},n_{1}\to\infty$ if $x_{(n_{0},n_{1})}/a_{(n_{0},n_{1})}$ is stochastically bounded, and write $x_{(n_{0},n_{1})}=o_{P}(1)$ if $x_{(n_{0},n_{1})}\to 0$ in probability. We write $x_{(n_{0},n_{1})}=\Omega_{P}(a_{(n_{0},n_{1})})$ if $x_{(n_{0},n_{1})}/a_{(n_{0},n_{1})}$ is stochastically bounded away from zero, and $x_{(n_{0},n_{1})}=\Theta_{P}(a_{(n_{0},n_{1})})$ if both $O_{P}(\cdot)$ and $\Omega_{P}(\cdot)$ hold.
@@
We consider binary classification under class imbalance, where $y=0$ denotes the majority class and $y=1$ denotes the minority class. Let the class-conditional distributions be $\bm{x}\mid(y=0)\sim\mathcal{P}_{0},~\bm{x}\mid(y=1)\sim\mathcal{P}_{1}.$ We observe $n_{0}$ majority samples and $n_{1}$ minority samples $(n_{0}>n_{1})$, and we may augment the data with $\tilde{n}$ synthetic minority samples. Synthetic covariates are drawn from a generator-induced distribution $\mathcal{P}_{\text{syn}}$; since the generator is typically trained on the observed data, $\mathcal{P}_{\text{syn}}$ may depend on $(n_{0},n_{1})$ and should therefore be written more precisely as $\mathcal{P}_{\text{syn},n_{0},n_{1}}$. We suppress this dependence for notational simplicity.
Let $\ell(\bm{\theta};\bm{x},y)$ be a loss function with parameter $\bm{\theta}\in\Theta$. Common choices include cross-entropy/logistic loss, squared loss, hinge loss, and exponential loss (see Section 10 of the Supplementary Materials for detailed discussions). To treat the two classes symmetrically at the population level, we define the balanced population risk
@@
and suppose there exists a minimizer $\bm{\theta}^{*}=\mathop{\rm arg\min}_{\bm{\theta}\in\Theta}\mathcal{R}(\bm{\theta}).$ Augmenting the training data with $\tilde{n}$ synthetic minority samples changes the effective class proportions (see Section 10 of the Supplement for a discussion of specific synthetic data generators used to produce these samples). Let $\pi_{0}=\frac{n_{0}}{n_{0}+n_{1}+\tilde{n}},~\pi_{1}=\frac{n_{1}}{n_{0}+n_{1}+\tilde{n}},~\tilde{\pi}=\frac{\tilde{n}}{n_{0}+n_{1}+\tilde{n}}$; so that $\pi_{0}+\pi_{1}+\tilde{\pi}=1$. The resulting synthetic population risk is
@@
with minimizer $\tilde{\bm{\theta}}=\mathop{\rm arg\min}_{\bm{\theta}\in\Theta}\tilde{\mathcal{R}}(\bm{\theta})$. In practice, we observe $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n_{0}+n_{1}}$ and generate synthetic samples $\{\tilde{\bm{x}}_{i}\}_{i=1}^{\tilde{n}}\sim\mathcal{P}_{\text{syn}},~\tilde{y}_{i}\equiv 1.$ We then minimize the empirical augmented risk
@@
where $\{\bm{x}_{i}^{(0)}\}_{i=1}^{n_{0}}$ and $\{\bm{x}_{i}^{(1)}\}_{i=1}^{n_{1}}$ are the observed majority and minority samples. Let $\hat{\bm{\theta}}=\mathop{\rm arg\min}_{\bm{\theta}\in\Theta}\hat{\mathcal{R}}(\bm{\theta}).$ Our goal is to estimate the balanced-risk optimizer $\bm{\theta}^{*}$ from imbalanced data, and to understand how synthetic augmentation, through $\tilde{n}$ and $\mathcal{P}_{\text{syn}}$, affects the balanced excess risk $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$.